the position of a train along it tracks as a function of time is plotted as a graph. at t1=6.0secs x1=48.5m at t2=8.50secs x2=73.5m. find the average velocity of this body in k/h.
1
Expert's answer
2013-03-05T09:12:13-0500
V - an average speed
S - a total distance
t - a total time
V = S/t
S = x2 - x1
t = t2 - t1.
Therefore:
V = (x2 - x1)/(t2 - t1) = (73.5m-48.5m)/(8.5sec-6sec) = 25m/2.5sec = 10m/sec
In km/h:
V = 10m/sec * 3600sec/1h * 1km/1000m = 10*3.6 km/h = 36 km/h